I thought I submitted this yesterday, but I did not get any answers on it. I need to see if I came up with the correct answers. The * sign is used as the mulitplication sign in this problem.

Problem #1)
x-6/x-3 = 2x/x-3
x-3/1 (x-6/x-3) = x-3/1 (2x/x-3)
x(x-3/1) - x-3/1 * 6/x-3 = 2x
x^2 - 3x-6 = 2x
-2x -2x
x^2 - 5x-6 = 0
(x+1)(x-6) = 0
x=-3 x=6

Problem #2 Word problem
In a city of 25,000 homes, a survey was taken to determine the number with cable television. Of the 300 homes survey, 210 had cable television. Estimate the number of homes in the city that had cable television.
210/300 = x/25,000
25,000 * 210/300 = x/25,000 * 25,000
17,500 = x
Thanks for you help

Answer by ankor@dixie-net.com(12701)

Problem #1)
x-6/x-3 = 2x/x-3
You don't have to do all that, multiply both sides by (x-3) and you have:
x - 6 = 2x
-6 = 2x-x
x = -6
:
:
Problem #2 Word problem
In a city of 25,000 homes, a survey was taken to determine the number with cable television. Of the 300 homes survey, 210 had cable television. Estimate the number of homes in the city that had cable television.
210/300 = x/25,000
25,000 * 210/300 = x/25,000 * 25,000
17,500 = x
:
This is right